OpenStudy (ledah):

What is the slope of the line through the points (-2, 7) and (3, -3)?

OpenStudy (p0sitr0n):

\[y=mx+b, where m =\Delta y/ \Delta x, \Delta y = y2 -y1, \Delta x = x2 -x1\] so, -3 - 7 / 3- (-2) = -10 / 5= -2

OpenStudy (ledah):

Nevermind Its -2 I got it. just miscalculations